If a house had 20 windows and ony a quarter used to be repaired,then how many need to be repaired?
1
Expert's answer
2019-11-25T12:08:25-0500

Since a quarter of 20 windows equals five (20*1/4 =20/4=5), then 5 windows should be repaired.

Answer: 5 windows.